how many ounces of a 2% cleaning solution should be mixed with a 5% cleaning solution to obtain 20 ounces of a 3.8% cleaning solution?
Answer by Alan3354(69443)

Same problem, different numbers.
---
A dairy needs 235 gallons of milk containing 5% butterfat. How many gallons each of milk containing 7% butterfat and milk containing 2% butterfat must be used to obtain the desired 235 gallons?
---------------------------
Let n be the amount of 7% butterfat milk. Then:
.07n+.02(235-n)=.05(235)
7n+2(235-n)=5(235)
5n+470=1175
5n=705
n=141
The dairy needs 141 gallons of 7% butterfat milk, and 94 gallons of 2% butterfat milk for its desired mixture.